LOG-IN
Displaying report 1-1 of 1.
MIF (ITF Control)
Print this report.
yutaro.enomoto - 20:19, Sunday 02 December 2018 (7239)Get code to link to this report
X Arm Cavity has been Locked by the Main Laser PDH Loop

[Nakano, Miyazaki, T Yamamoto, Enomoto]

X arm cavity has been locked to the main laser!!

That means we achieved the hand-over from ALS CARM to (main laser) CARM loop, and X arm has been kept locked only using the PDH signal taken at REFL port.
The following figures show the IR transmission of X arm, and OLTF of CARM loop we took.

As you can see from the first figure, IR kept resonating in X arm even after ALS CARM loop was disabled (IN2_EN set 0).
The settings of CARM servo at this moment is:
IN1 pol: -
IN1 gain: 31dB (26dB was also OK)
FAST gain: 12dB
# of Common Boosts: 2
All others disabled
MC servo IN2 gain: -30dB
SR 560: gain 5, 30Hz LPF, pol +

We succeded the hand-over several times, and the procedure of the hand-over is the following:
- turn off many instruments to reduce noise in ALS (details to be reported)
- request ALS_PDHX to be ALS_CARM_LOCKED_WITH_HIGH_BANDWIDTH
- adjust the offset frequency of PLL so that TEM00 mode of main laser is at the resonance (usually the offset is around 40.027 MHz or 40.052 MHz)
- enable IN1 of CARM servo with IN gain of 31dB (26dB was also OK)
- gradually decrease the IN2 gain from 14dB to -32dB (14dB --> 6dB --> 4dB --> 0dB --> -32dB)
- disable IN2
- adjust offset of CARM servo if needed.

It can be said that this is the completion of frequency stablization system of KAGRA, or the achieve of lock of the first degree of freedom of MIF.

Images attached to this report
7239_20181202115531_firsthandover2ircarm.png 7239_20181202115554_mokubodeanalyzerdata20181202183149screenshot.png
Comments related to this report
takashi.uchiyama - 20:26, Sunday 02 December 2018 (7240)
Great achievement!!

Thank you.
shinji.miyoki - 23:18, Sunday 02 December 2018 (7242)
So grate!
masayuki.nakano - 12:30, Monday 03 December 2018 (7255)

The attached file is the configuration of the PSL room, the compressor, the air conditioner, and the FFU units around the PR, the MC, the BS, and the SR area during the CARM locking with IR. 

yutaro.enomoto - 12:51, Monday 03 December 2018 (7256)

Here is the out of loop stability of ALS, estimated with X-arm-stabilized main laser frequency.
RMS is taken below 59 Hz to exclude the effect of line noise, and X arm cavity pole of ~ 30 Hz is included.
The RMS between 0.1Hz and 59Hz was as low as 13 Hz.